Year A B 1 10% 9% 2 -15% 16% 3 14% -10% 4 13% -12% 5...

Year

A

B

1

10%

9%

2

-15%

16%

3

14%

-10%

4

13%

-12%

5

10%

21%

6

13%

11%

7

-12%

14%

8

10%

-10%

9

10%

-12%

10

-19%

21%

What is the correlation and covariance between these two assets?
Answer #1
A D = Return-Average Return D^D B D = Return-Average Return D^D D(A)*D(B)
0.1 0.066 0.004356 0.09 0.042 0.001764 0.002772
-0.15 -0.184 0.033856 0.16 0.112 0.012544 -0.02061
0.14 0.106 0.011236 -0.1 -0.148 0.021904 -0.01569
0.13 0.096 0.009216 -0.12 -0.168 0.028224 -0.01613
0.1 0.066 0.004356 0.21 0.162 0.026244 0.010692
0.13 0.096 0.009216 0.11 0.062 0.003844 0.005952
-0.12 -0.154 0.023716 0.14 0.092 0.008464 -0.01417
0.1 0.066 0.004356 -0.1 -0.148 0.021904 -0.00977
0.1 0.066 0.004356 -0.12 -0.168 0.028224 -0.01109
-0.19 -0.224 0.050176 0.21 0.162 0.026244 -0.03629
Sum = 0.34 0.15484 Sum = 0.48 0.17936 Sum = -0.10432
n =
[Number of Observations]
10 10 n =
[Number of Observations]
10 10 n =
[Number of Observations]
10
Average Return =
[Sum of Returns/n]
0.034 Variance =
[Sum of D^D/n]
0.015484 Average Return =
[Sum of Returns/n]
0.048 Variance =
[Sum of D^D/n]
0.017936 Covariance =
[Sum of D(A)*D(B)/n]
-0.01043
Standard Deviation =
[Square Root of Variance]
0.124435 Standard Deviation =
[Square Root of Variance]
0.133925

Correlation = Covariance/(Std Dev of A*Std Dev of B) = -0.01043/(0.124435*0.133925) = -0.62586